Shift summary, Eve 08/02/99

Patrick Toale (toale@cuhep.Colorado.EDU)
Mon, 2 Aug 1999 23:02:49 -0600 (MDT)

Shift Summary for Eve shift, 08/02/99

On Shift: Pat T., Hogan N.

Hours of beam delivered at all intensities: 8.5/8.5
Hours of beam delivered at nominal intensity: 8.5/8.5
Hours of E832 data (special runs not included): 7.8/8.5
Hours of special runs (laser scans, muon runs...): 0.1/8.5

Accounting of lost time: new tapes, muon run, polarity flip

========================================================================
Runs ended this shift: 14090, 14091
Run 14090 - BEAM_KE3HOG
Run 14081 - MUON

Tapes brought to FCC this shift: QKE001-QKE009, QKB072

========================================================================
Accesses during shift: none

========================================================================
Description of Shift: smooth sailing

run spill time comment
------------------------------------------------------------------------
14090 15:30 Inherit smooth run.

21:45 Tapes full, stop run 14090.
Onsplit requests tapes QKPE09 and QKCC09,
we obey.

14091 22:00 Start MUON run 14091.

22:07 Stop run 14091, ramp magnet up with negative
polarity.

14092 22:32 Start BEAM_KE3HOG run 14092. Get crate 53 DYC full,
a few kfrends fixes it.

24:00 Hand-off to owl shift.

========================================================================
Follow up for next shift:

Previous estimate for the length of run necessary to gather a sufficient
number of Ke3's was overestimated by a factor of 3 or so. So the tentative
plan is to take at least one more tape set of Ke3's and then go back to
increased intensity. NOTE: Magnet should be set to positive polarity
before going to higher intensity running.

Evening shift was told to label many ke3hog tapes, too many it seems.

Thanks,
Pat Toale
--------------------------------------------------------------
Graduate Research Assistant
Department of Physics University of Colorado at Boulder
Office: F-231 Phone: (303) 492-4363
E-Mail: toale@cuhep.colorado.edu